OpenStudy (anonymous):

i need two equivalent fractioin for 2/10

OpenStudy (anonymous):

\[\LARGE \frac{2\div 2}{10\div 2}=answer1\] \[\LARGE \frac{2\times 2}{10\times 2}=answer 2\]

OpenStudy (anonymous):

thanks i also need two for each of this one 1/6, 1/4, 5/9, 1/2, 3/11, 6/12, 11/13, 2/3

OpenStudy (anonymous):

Double and halve them as well